respected sir,
1] Area of a right-angled triangle is doubled while the base is reduced to its half. What % increase occurs in its height?
plz help me to solve this geometry problem...
thank u.

Answer by Alan3354(69443)

Area of a right-angled triangle is doubled while the base is reduced to its half. What % increase occurs in its height?
------------
A = bh/2
If b is half, then h has to be 4x to double the area.
A = b*h/2
2A = (b/2)*h*x/2 = (x/2)*bh/2
x = 4A/A
x = 4
--------
The factor of increase is 4.
h increased from 1h to 4h, an increase of 3h --? 300% NOT 400%
4 times as much is a 300% increase.
Despite what dufuses on TV say, it is NOT a 400% increase.